OpenStudy (anonymous):

I WILL MEDAL AND FAN Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

Ill give you the expression

OpenStudy (anonymous):

Expression #1(3d)3(d) Expression #2 (3d2)2

OpenStudy (anonymous):

@Helpmeplease99

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Ell

OpenStudy (ell):

I'm really sorry, im not the best in math. Hope you find the right help though! God bless! :)

OpenStudy (anonymous):

:) its ok thank you you too

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@countonme123

OpenStudy (anonymous):

Aww thanks ell

OpenStudy (ell):

np:) Hey, @timo86m might be able to help:)

OpenStudy (anonymous):

Thank you :))

OpenStudy (ell):

np:)

OpenStudy (unklerhaukus):

i dont understand the notation you have used @HelloGoodmorning

OpenStudy (anonymous):

notation do you mean the exponents?

OpenStudy (unklerhaukus):

are these the expressions? Expression #1 \((3d)^3(d)\) Expression #2 \((3d^2)^2\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

can you simplify the first expression? ( use \((ab)^n=a^nb^n\) )

OpenStudy (anonymous):

is the b turning into a one like 3*1^3=3^3*1^3 which equals 18*1=18

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[(3d)^3(d)=3^3d^3\times d=\]

OpenStudy (unklerhaukus):

\[3^3=3\times3\times3=\]

OpenStudy (anonymous):

27

OpenStudy (unklerhaukus):

right so \[(3d)^3(d)=27d^3\times d=\] ___ now simplify the d's use \(a=a^1\) and \(a^n\times a^m=a^{n+m}\)

OpenStudy (anonymous):

? confused @UnkleRhaukus

OpenStudy (unklerhaukus):

well in you case a=d , n=3 and m=1

OpenStudy (anonymous):

Im really confused for solving it

OpenStudy (unklerhaukus):

i am not sure where you are confused?

OpenStudy (anonymous):

Im confused because you've asked me to simplify the d

OpenStudy (unklerhaukus):

what is \(d^3\times d=(d\times d\times d)\times d=?\)

OpenStudy (anonymous):

d^4

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

ok so what do you get for the most simplified form of expression #1?

OpenStudy (anonymous):

27^4

OpenStudy (anonymous):

But the question is only asking for which one has more exponents You dont have to solve

OpenStudy (unklerhaukus):

i hope you mean 27d^4 now simplify the second expression (it is much easier to compare the expression when that are as simplified as possible )

OpenStudy (anonymous):

Okay the second one is 3d^4

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

nearly, try that one again \[(3d^2)^2=3^2d^{2\times2}=\]

OpenStudy (anonymous):

ok so that equals 3d^12 or 3d^27

OpenStudy (unklerhaukus):

3^2=?

OpenStudy (anonymous):

3*3=9

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

so the second answer has more exponents??

OpenStudy (unklerhaukus):

what was you final simplified form of expression#2 ?

OpenStudy (anonymous):

3d^27

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

How come i thought i was supposed to do 3*3*3

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[(3d^2)^2=3^2d^{2\times2}=3\times3\times d^4=\]

OpenStudy (anonymous):

So the first expression has more?

OpenStudy (anonymous):

@UnkleRhaukus @Darrius @Helpmeplease99